OpenStudy (anonymous):

assume that you randomly select 4 cards from a deck of 52. what is the probability that all the cards selected are spades?

OpenStudy (anonymous):

13

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

wait, do you replace teh cards each times?

OpenStudy (anonymous):

its is not 13

OpenStudy (anonymous):

13 is not a probability

OpenStudy (anonymous):

9/13

OpenStudy (anonymous):

you need to know if you replace the cards: Independent or don't replace: dependent

OpenStudy (anonymous):

it doesnt say you do so i assume you dont

OpenStudy (anonymous):

its 9/13 try that

OpenStudy (anonymous):

that is wrong i tried it

OpenStudy (anonymous):

then you didn't write the question correctly?

OpenStudy (anonymous):

Assume that you randomly select 4 cards from a deck of 52. What is the probability that all of the cards selected are spades?

OpenStudy (anonymous):

I don't see how 9/13 is wrong but oh well

OpenStudy (anonymous):

in cards of total 52 there are 13 hearts 13 diamonds 13 clubs and 13 spades... the probability of picking up a spade is.... sum of spades/total cards= 13/52= 1/4

OpenStudy (anonymous):

Yes 1 spade but he said 4

OpenStudy (anonymous):

u take 4 spades then its 13C4/52C4

OpenStudy (anonymous):

that did not work either

OpenStudy (anonymous):

Okay can you write the whole question, because it changes the way you work it

OpenStudy (anonymous):

Assume that you randomly select 4 cards from a deck of 52. What is the probability that all of the cards selected are spades? thats the whole question